If Steven can mix 20 drinks in 15 minutes, Sue can mix 20 drinks in 10 minutes, and Jack can mix 20 drinks in 15 minutes, how much time will it take all 3 of them working together to mix the 20 drinks?
A= 2 minutes and 44 seconds
B= 2 minutes and 58 seconds
C= 2 minutes and 10 seconds
D= 3 minutes and 26 seconds
E= 4 minutes and 15 seconds

Answer by jorel1380(3719)

20/15+20/10+20/15=20/x
40/30+60/30+40/30=20/x
140/30=20/x
600=140x
60/14=x=30/7 mins..